A jeweler has 2 rings, each weighing 24 g, made of an alloy of 5% silver and 95% gold. She decides to melt down the rings and add enough silver to reduce the gold content to 80%. How much silver should she add?
Answer by stanbon(75887)

A jeweler has 2 rings, each weighing 24 g, made of an alloy of 5% silver and 95% gold. She decides to melt down the rings and add enough silver to reduce the gold content to 80%. How much silver should she add?
--------------------
Equation:
silver + silver = silver
0.05*48 + 1.00s = 0.20(48+s)
Multiply thru by 100 to get:
5*48 + 100s = 20*48+20s
80s = 15*48
s = 9 grams (amt. of pure silver she needs to add to the mix)
=======================
Cheers,
Stan H.